OpenStudy (anonymous):

Okay guys, last question. I promise, I am just studying for a Math test that is coming up. Solve the following equation: -4y+8=4(2y-2)-2(-16+8y) A) y=-4 B) y=4 C) y=-2 D) y=-6

OpenStudy (ranga):

Use Distributive Property to get rid of all the parenthesis on the right first. a(b + c) = ab + ac. Do the step and I will help you with the next step.

OpenStudy (anonymous):

Okay, I understand what the distributive property is. The only problem is I don't see where I can start with that in this equation. Do I have to do what is inside the ( ) first?

OpenStudy (ranga):

You have 4(2y-2). You can use distributive property there. You also have -2(-16+8y) where you can use distributive property again.

OpenStudy (anonymous):

Okay, so should I use distributive property on both of them like this? -4y+8=8y-8-32-16y Is that right?

OpenStudy (ranga):

You are almost correct. Check the sign of 32. Should it be -32 or +32?

OpenStudy (anonymous):

+32 I believe.

OpenStudy (ranga):

Correct. So you have -4y + 8 = 8y - 8 + 32 - 16y Next bring all the y terms to the left of the equal sign and take all the constant terms to the right of the equal sign. Note: Switch signs when you take a left hand side term to the right or when you take a right hand side term to the left. For example, 5x - 2 = 3x + 6 can be written as 5x - 3x = 6 + 2. Do the same for the above equation.

OpenStudy (anonymous):

So are you saying that I should switch the equation vise versa before I do this part?

OpenStudy (ranga):

We want to solve for y. So we are gathering up all the y terms on the left and all the constant terms on the right.

OpenStudy (anonymous):

-4y+8=8y-8+32-16y -4y+8=8y+32 -8 -4y=8y+32 Right so far? Or no?

OpenStudy (ranga):

Not correct. Okay it looks like you are used to doing it the long way. First I will take an example and do it the long way. Then I will do the same the easier way. Example: 5x - 2 = 3x +6 To get rid of the -2 on the left, add +2 to both sides: 5x -2 = 3x + 6 +2 = + 2 5x = 3x + 8 To get rid of 3x on the right subtract 3x from both sides: 5x = 3x + 8 -3x -3x 2x = 8 x = 4.

OpenStudy (ranga):

Now I will show you the easy way to solve 5x - 2 = 3x + 6 Bring the x terms to the left and constant terms to the right. When you switch sides, simply remember to change the sign. So a 3x on the right when taken to the left will become -3x and a -2 on the left when taken to the right will become +2. So 5x - 3x = 6 + 2 2x = 8 x = 4.

OpenStudy (anonymous):

Hmmmm, I do admit that I am used to doing it the long way. I am not familiar with doing it the easy way.

OpenStudy (ranga):

Okay, if you stll prefer to do it the long way then proceed with where we left off: -4y+8=8y-8+32-16y On the right you can simplify first by adding or subtracting the like terms

OpenStudy (anonymous):

Okay, I prefer to do that because nobody really ever showed me the easy way. If they did, I must have not been paying attention. Anyways.. -4y+8=8y-8+32 So that is done.

OpenStudy (ranga):

No. not yet done. You have to watch for the signs carefully. On the right we had 8y-16y = ? Also on the right we have -8 + 32 = ?

OpenStudy (anonymous):

Okay, well -8+32=24 and wouldn't 8y-16y=-8y So then wouldn't we be at -4y+8=-8y+24?

OpenStudy (ranga):

Excellent! Now do your long way and solve for y.

OpenStudy (ranga):

Subtract 8 from both sides of the equation first.

OpenStudy (anonymous):

-4y+8=-8y+24 -8 -8 -4y=-8y+16 Am I still on track?

OpenStudy (ranga):

Yes! What should be the next step?

OpenStudy (anonymous):

Well don't we still have to isolate y?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

Okay, so that is the next step I do believe. The only problem is that there is still a y on both sides of the equation. That does not seem right.

OpenStudy (ranga):

Add 8y to both sides to get rid of the y term on the right.

OpenStudy (anonymous):

Okay, so like this? -4y=-8y+16 +8y +8y 4y=16 4y/4=4 16/4=4 I think I got the answer, the answer is 4! :D

OpenStudy (ranga):

Excellent!

OpenStudy (ranga):

But if you learn to do it the easy way the whole problem can be done in just 5 lines. It will save you a lot of time doing homework and taking tests. So here it is: -4y+8=4(2y-2)-2(-16+8y) -4y + 8 = 8y -8 + 32 - 16y -4y - 8y + 16y = -8 - 8 + 32 4y = 16 y = 4 Good luck!

OpenStudy (anonymous):

Okay, good stuff. I will have to study the easy way now, I just have been so used to doing these problems the long way. Thanks for actually spending all of this time on helping me out, I owe you a huge thank you sir. Anyways, thank you and have a good night!

OpenStudy (ranga):

You are welcome. Good night.